OpenStudy (jenniferjuice):

Simplify completely quantity 3 x plus 18 over 18. x 3x quantity x plus 6 over 6 quantity x plus 18 over 6

OpenStudy (imstuck):

\[\frac{ 3x+18 }{ 18 }\]

OpenStudy (imstuck):

\[\frac{ 3(x+6) }{ 18 }\]

OpenStudy (imstuck):

now reduce between the 3 and the 18 to get\[\frac{ (x+6) }{ 6 }\]and that's simplified completely. Factor first and you're good to go!

OpenStudy (imstuck):

You got that, right? As long as you understand it...

OpenStudy (imstuck):

TY for the medal!

OpenStudy (jenniferjuice):

yess thank youuu